# Explicit conserved operators for a class of integrable bosonic networks from the classical Yang-Baxter equation

### v1

@@ 4 An example
For the case $m=2$, choosing
$\mathcal{B}=-\frac{1}{2}\left(\begin{array}[]{cc}\displaystyle{K}_{1}&\displaystyle{J}\\ \displaystyle{J}&\displaystyle{K_{2}}\end{array}\right)\qquad\qquad J,\,K_{1},\,K_{2}\in\,{\mathbb{R}}$
@@
$\mathcal{H}=\mathcal{U}(N^{2}_{a}+N^{2}_{b}-I)$ $-\frac{K_{1}}{2}(a_{1}^{\dagger}b_{1}+b_{1}^{\dagger}a_{1})-\frac{K_{2}}{2}(a_{2}^{\dagger}b_{2}+b_{2}^{\dagger}a_{2})-\frac{J}{2}(a_{2}^{\dagger}b_{1}+b_{2}^{\dagger}a_{1}+a_{1}^{\dagger}b_{2}+b_{1}^{\dagger}a_{2}).$
Making a canonical transformation to boson operators $\{c_{j},\,c_{j}^{\dagger}:j=1,2,3,4\}$ through
@@ 5 Discussion
The main objective of this work was to provide explicit forms for the conserved operators of the Hamiltonian (6). These conserved operators are expressed through (43) and (44). They were obtained by quantisation of classical counterparts obtained through a Poisson algebra approach. These results confirm that the models studied in [73] associated with complete bipartite graphs generalise to models associated with arbitrary bipartite graphs.
@@ Acknowledgments
This research was supported by the Australian Research Council through Discovery Project DP200101339, Quantum control designed from broken integrability. We thank Angela Foerster for comments and correspondence. We acknowledge the traditional owners of the land on which The University of Queensland (St. Lucia campus) operates.

### v2

@@ 4 An example
To illustrate the results, we work through the minimal non-trivial example for which $m=2$. This correspond to the choice $K_{2,2}$ for the bipartite graph. The most general symmetric matrix that can be used to weight the edges of the graph has three independent parameters, as depicted in Fig. 1. For this parametrisation
$\mathcal{B}=-\frac{1}{2}\left(\begin{array}[]{cc}\displaystyle{K}_{1}&\displaystyle{J}\\ \displaystyle{J}&\displaystyle{K_{2}}\end{array}\right),\qquad\qquad J,\,K_{1},\,K_{2}\in\,{\mathbb{R}}$, (48)
@@
$\mathcal{H}=\mathcal{U}(N^{2}_{a}+N^{2}_{b}-I)$ $-\frac{K_{1}}{2}(a_{1}^{\dagger}b_{1}+b_{1}^{\dagger}a_{1})-\frac{K_{2}}{2}(a_{2}^{\dagger}b_{2}+b_{2}^{\dagger}a_{2})-\frac{J}{2}(a_{2}^{\dagger}b_{1}+b_{2}^{\dagger}a_{1}+a_{1}^{\dagger}b_{2}+b_{1}^{\dagger}a_{2}).$ (49)
[…]
Figure 1: Graphical presentation of the complete bipartitie graph $K_{2,2}$. The vertex sets coloured blue and teal are assigned labels 1 and 2. Weights are assigned to the edges such that the associated matrix $\mathcal{B}$ given by (48) is symmetric, leading to the Hamiltonian (49).
Making a transformation to boson operators $\{c_{j},\,c_{j}^{\dagger}:j=1,2,3,4\}$ through
@@ 5 Discussion
The main objective of this work was to provide explicit forms for the conserved operators of the Hamiltonian (6). These conserved operators are expressed through (44) and (45). They were obtained by quantisation of classical counterparts obtained through a Poisson algebra approach. These results confirm that the models studied in [77] associated with complete bipartite graphs generalise to models associated with arbitrary bipartite graphs. We remark that this generalised class of bosonic models is distinct from those derived in other studies [54, 55, 56] that are also founded on solutions of the classical Yang-Baxter equation.
@@ Acknowledgments
This research was supported by the Australian Research Council through Discovery Project DP200101339, Quantum control designed from broken integrability. We thank Angela Foerster for comments and correspondence, and Taras Skrypnyk for helpful feedback. We acknowledge the traditional owners of the land on which The University of Queensland (St. Lucia campus) operates.